OpenStudy (anonymous):

add (7+3i)+(-9-8i)=?

OpenStudy (anonymous):

you add normally (7+3i)+(-9-8i)=(7-9)+(3i-8i)=-2-5i

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem